A heart typically beats 4320 times in an hour. At this rate, how many times does a heart beat in 8 minutes?
Answer by jim_thompson5910(35256)


4320/60 = x/8 ... Note: 1 hour = 6 min

4320*8 = 60x

34560 = 60x

34560/60 = x

576 = x

x = 576

So a heart typically beats 576 times in 8 minutes.